OpenStudy (anonymous):

@Mathatics hi is1/2 less or more than 1/5 @Mathatics hi is1/2 less or more than 1/5 @Mathematics

OpenStudy (anonymous):

Simplify the expression -2x2(4x − 3).

OpenStudy (anonymous):

it is more. \[\frac{1}{2}=\frac{5}{10}\] \[\frac{1}{5}=\frac{2}{10}\]

OpenStudy (anonymous):

Thanx